A cylindrical rod of length 2.0 m and cross-sectional area 8.0 × 10–5 m2 is made of hard steel. The rod is placed under a tensile stress of 200 MPa.

Which one of the following is the best estimate of the force that must be applied to this hard steel rod to provide this stress?


1
Expert's answer
2021-01-29T20:01:47-0500

By definition, the stress is:


"\sigma = \dfrac{F}{S}"

where "F" is the force that must be applied to this hard steel rod, and "S = 8\times 10^{-5}m^2" is its cross-sectional area. Since "\sigma = 200\times 10^6Pa", the force is:


"F = \sigma S\\
F = 200\times 10^6\cdot 8\times 10^{-5} = 16000N"

Answer. 16000 N.